Shift report 20011023 16:00-24:00

From: Apache (apache@pii3.brahms.bnl.gov)
Date: Tue Oct 23 2001 - 23:47:28 EDT

  • Next message: Apache: "Shift report 20011024 00:00-08:00"

    ------------------------------------------------------------
    Date: Tue Oct 23 23:47:27 2001, Shift: 16:00-24:00
    Supervisor: JC, with: RW
    Log book#10 pages 53 - 53
    
    Shift Summary:
    ------------------------------------------------------------
    1. No beam at all. A few attempts at ramping the beam failed.
    Message "Rhic Quench link interlock problems" constantly flashing.
    The last message at 00:00 : "Rhic controls problem. An expert 
    is looking into it"
    
    2. No data runs.
    DVM and pedestals runs - 5497-5499.
    
    3. No detector problems noticed.
    ------------------------------------------------------------
    



    This archive was generated by hypermail 2b30 : Tue Oct 23 2001 - 23:48:07 EDT